tonys mother is 8 more than twice his age.his father is three years older than his mother is. if all three family members have lived 94 years, how old is each family member
Answer by dkppathak(439)

let tonny age is x yrs
as per given conditions
mother age is 2x+8
father age is 3 yrs amore than mother
father age is 2x+8+3
2x+11
x+2x+8+2x+11=94
5x+19=94
5x=94-19
5x=75
x=75/5=15
tonny is 15 yrs
mother is 38 yrs
father is 41 yrs